Write nernst equation for Zn and Cu electrode considering both behave as metal metal ion electrode
1
Expert's answer
2020-10-26T14:54:12-0400

The overall reaction is:

Zn(s) + Cu2+(aq) --> Zn2+(aq) + Cu(s)

At 298K, the Nernst equation in general can be represented as


"E_{cell}=E^0_{cell}-\frac{0.0591V}{n}logQ" .

In this case "n=2" as there are 2 electrons transferred in the cell reaction, and

"Q=\frac{[Zn^{2+}]}{[Cu^{2+}]}". Therefore, the overall equation is:


"E_{cell}=E^0_{cell}-\frac{0.0591V}{2}log\frac{[Zn^{2+}]}{[Cu^{2+}]}"